Two firms with a constant MC and no fixed cost compete in a Cournot fashion. Firm A's MC=20 and Firm B's MC=40.

Market demand is given by x=180-p. The government imposes a per unit tax of ?=10 which the producer pays. What is market price?
Answer #1

MC = 20 MC₂ = 40 x=180-p = p = 180-X tax ; t = 10 - 160 - x + lo P = 190-X Profit function P= 190 - x, ~ X2 ^, = P,X, - MC, XIX, & X2 = 170 X, + 2x2 = 100 x 2 2x + x2 = 1 to 2x, & 4X₂ i 3n - 3X2 = - 130 X 2 = 130 2x, & 130 - 170 8x = 120 - 130 2x, =